OpenStudy (anonymous):

Find the slope of the line passing through the points (-3, -10) and (4, 5).

OpenStudy (anonymous):

Slope=y2-y1/x2-x1 Now, plug in the numbers.

OpenStudy (anonymous):

slope=change in y/change in x=(y2-y1)/(x2-x1) and here y2=5, y1=-10, x2=4, and x1=-3. We have: slope=(5-(-10))/(4-(-3)= 15/7

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ -10}})\quad &({\color{red}{ 4}}\quad ,&{\color{blue}{ 5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)